# Dynamic Investment-Driven Insurance Pricing: Equilibrium Analysis and Welfare Implication

### v1

@@ 2.3 Optimization Objective
Definition 1
For each $t\in[0,T]$, $a\triangleq\Big{\{}a_{s}=(x_{s},y_{s}):t\leq s\leq T\Big{\}}$ is an admissible strategy for the insurer with initial wealth $m_{t}=m$ and denoted as $a\in\mathcal{A}_{t}$ if it satisfies:
@@
    For any $(t,m)\in[0,T]\times\mathbb{R}$, $\mathbb{E}_{t,m}\Big{[}\sup\limits_{t\leq s\leq T}\left|m_{s}\right|^{2}\Big{]}<\infty$;
@@
$V(t,m)\triangleq\sup_{a\in\mathcal{A}_{t}}\mathbb{E}_{t,x}\Big{[}U(m(T))\Big{]},$ (2.2)
where $V(\cdot,\cdot)$ denotes the value function, with the boundary condition $V(T,m)=U(m)$. Based on this optimization problem, we now provide the definition of the general equilibrium of the insurance market.
Definition 2
A stationary Markovian competitive equilibrium consists of: an insurance price process $\Theta$, an insurance demand process $D$, an insurer’s wealth process $M$, an insurance supply process $X$, and an insurer’s investment process $Y$, that are compatible with the insurer’s optimization objective (2.2) and the market clearing condition $D=X$.
@@ 3.1 Solution to Decision Problem
$V_{t}+\sup_{x\geq 0,y}\Bigg{\{}V_{m}\Big{(}x\theta l+y(\mu-r)+mr\Big{)}+\frac{1}{2}V_{mm}\Big{(}x^{2}\eta^{2}+2\rho x\eta y\sigma+y^{2}\sigma^{2}\Big{)}\Bigg{\}}=0.$ (3.1)
@@
For $\lambda^{\ast}\geq 0$, it requires that $\rho\geq\frac{\theta l\sigma}{(\mu-r)\eta}$. Substituting this result back into the HJB equation, we obtain
@@
For $x^{\ast}>0$, we require $\rho<\frac{\theta l\sigma}{(\mu-r)\eta}$. Replacing them back into the HJB equation gives
@@
Proposition 1
Let $f(t,m,x,y)=V_{m}\Big{(}x\theta l+y(\mu-r)+mr\Big{)}+\frac{1}{2}V_{mm}\Big{(}x^{2}\eta^{2}+2\rho x\eta y\sigma+y^{2}\sigma^{2}\Big{)}$. Assume that $V_{m}>0$ and $V_{mm}<0$. $(x^{\ast},y^{\ast})$ is defined as follows:
@@
Then $f(t,m,x^{\ast},y^{\ast})=\sup\limits_{x\geq 0,y}f(t,m,x,y)$. Consequently, if $a^{\ast}=\Big{\{}a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):(x^{\ast}_{s},y^{\ast}_{s})=(x^{\ast}(t,m^{\ast}_{s}),y^{\ast}(t,m^{\ast}_{s}):t\leq s\leq T\Big{\}}\in\mathcal{A}_{t}$ and the HJB equation (3.1) admits a solution $V$, then $a^{\ast}$ is the optimal control of problem (2.2), where $M^{\ast}=\{m^{\ast}_{s}:s\geq t\}$ is a unique solution of SDE (2.1) when the process $a$ is replaced with $a^{\ast}$.
@@ 3.2 Case of CARA Utility
Proposition 2
Given the price process $\Theta$, for an insurer with CARA utility and initial wealth $m$ at time $t$, the optimal underwriting and investment strategy is $a^{\ast}=\Big{\{}a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):t\leq s\leq T\Big{\}}$, and the corresponding value function $V(t,m)$ is
@@
Proof. Following Yang and Zhang, (2005), we guess that the value function takes the form of (3.2). It can be efficiently verified that $a^{\ast}$ is admissible and $V$ satisfies the HJB equation (3.1). Therefore, they solve Problem (2.2). $\square$
Note that now the price $\theta_{s}$ is assumed to be given, and each insurer in the competitive market acts as a price taker. Although the equilibrium price would be eventually determined through the interaction of market supply and demand, the supply function $x^{\ast}_{s}(\theta_{s})$, as shown in (3.3) and (3.4), is deterministic at each point in time, and the demand function $d(\theta_{s})$ is also deterministic. As a result, the equilibrium price is not stochastic and varies only with time $s$. Insurers hold deterministic beliefs about this equilibrium price process.
@@ 3.3 Market Equilibrium
Now we analyze the equilibrium conditions in the insurance market. In equilibrium, the competitive price of insurance must clear the market, meaning that the aggregate supply of insurance meets the aggregate demand. We should hold the relationship:
@@
Theorem 1 (Market Equilibrium)
@@
    $\Theta^{\ast}_{z}=\Big{\{}\theta^{\ast}_{s}=\overline{\theta}:t\leq s\leq T\Big{\}}.$
@@
    $X^{\ast}_{z}(\Theta^{\ast})=\Big{\{}x^{\ast}_{s}(\theta^{\ast}_{s})=0:t\leq s\leq T\Big{\}},$
@@
    then there exists a stopping time $\tau_{f}=\inf\limits_{s\geq t}\Big{\{}\underline{\theta}>\gamma(1-\rho^{2})\frac{\eta^{2}}{l}e^{r(T-s)}+\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}\Big{\}}$. When $\tau_{f}\leq s\leq T$, the market fails. Thus, there is no equilibrium in the insurance market.
Proof. Based on Proposition 2, when $\phi_{s}>\rho$, the insurer actively participates in both underwriting and investment. In this case, the supply function is given by $x^{\ast}_{s}(\theta_{s})=\frac{\theta_{s}l\sigma-\rho(\mu-r)\eta}{\gamma(1-\rho^{2})\eta^{2}\sigma}e^{-r(T-s)}$, which is an upward-sloping linear curve and has an intersection with the downward-sloping demand curve. But to further guarantee the intersection is an equilibrium, we need
@@
When $\phi_{s}\leq\rho$, the insurer stops underwriting, and $x^{\ast}_{s}=0$. The equilibrium price becomes $\theta^{\ast}_{s}=\overline{\theta}$, and it should satisfy $\phi_{s}(\theta^{\ast}_{s})\leq\rho$. Then Condition (3.6) is derived. Beyond these two cases, when $\underline{\theta}$ is set relatively high, the intersection would violate $\underline{\theta}\leq\theta^{\ast}_{s}$, meaning there is no feasible equilibrium. This results in the condition specified in (3.7). $\square$
@@ 3.5 Case of Negative Loading
Proposition 5
Given $\frac{\rho}{1-\rho^{2}}<-\gamma\frac{\sigma}{\mu-r}\frac{\eta\overline{\theta}}{\Delta\theta}$, and the positive insurance market condition (3.5), there exists a stopping time $\tau_{n}=\inf\limits_{s\geq t}\Big{\{}\frac{\rho}{1-\rho^{2}}<-\gamma\frac{\sigma}{\mu-r}\frac{\eta\overline{\theta}}{\Delta\theta}e^{r(T-s)}\Big{\}}$. When $\tau_{n}\leq s\leq T$, the equilibrium loading premium $\theta^{\ast}_{s}$ becomes negative.
@@ 3.6 Welfare Implication
Theorem 2 (Social Welfare)
@@ 3.7 Case of Perfect Correlation
$\Theta^{\ast}=\left\{\theta^{\ast}_{s}=\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}:t\leq s\leq T\right\}.$
@@
$X^{\ast}(\Theta^{\ast})=\left\{x^{\ast}_{s}(\theta^{\ast}_{s})=\frac{\overline{\theta}-\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}}{\Delta\theta}:t\leq s\leq T\right\},$
@@
$Y^{\ast}(\Theta^{\ast})=\left\{y^{\ast}_{s}(\theta^{\ast}_{s})=\frac{\frac{\eta^{2}}{\sigma}\left(\frac{\mu-r}{\sigma}-\rho\frac{l}{\eta}\overline{\theta}\right)+\frac{\mu-r}{\gamma\sigma^{2}}l\Delta\theta e^{-r(T-s)}}{l\Delta\theta}:t\leq s\leq T\right\}.$
@@ 4.1 Market Equilibrium
$\mathrm{d}m_{t}=\left(x_{t}\theta_{t}l+y_{t}(\mu-r)-\frac{1}{2}y_{t}^{2}\epsilon_{t}+m_{t}r\right)\mathrm{d}t+x_{t}\eta\mathrm{d}W^{I}_{t}+y_{t}\sigma\mathrm{d}W^{S}_{t},$
where $\epsilon_{t}\in[0,\overline{\epsilon}]$ denotes the intensity of regulatory frictions and consists of $\mathcal{E}\triangleq\Big{\{}\epsilon_{t}:t\geq 0\Big{\}}$. And $\overline{\epsilon}$ is an exogenously given bar. Using a similar approach to the derivation in Section 3, we obtain the following proposition describing the insurer’s optimal strategy.
Proposition 7
Given the price process $\Theta$ and the regulatory cost process $\mathcal{E}$, for an insurer with CARA utility and initial wealth $m$ at time $t$, the optimal underwriting and investment strategy is $a^{\ast}=\Big{\{}a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):t\leq s\leq T\Big{\}}$, and corresponding value function $V(t,m)$ is
@@
Theorem 3 (Market Equilibrium with Regulatory Friction)
@@ 4.2 Welfare Implication
Lastly, we examine how regulatory frictions impact the social welfare of the insurance market and explore the conditions for the regulator to set an optimal cost.
Theorem 4 (Social Welfare with Regulatory Friction)
@@
Proof. The social welfare can be easily calculated by $w_{s}^{\ast}=\frac{1}{2}x_{s}^{\ast}\left(\overline{\theta}-\frac{\rho(\mu-r)\eta}{\left(1+\frac{\epsilon_{s}}{\gamma\sigma^{2}}e^{-r(T-s)}\right)l\sigma}\right)$. Define $\alpha_{s}=\frac{1}{1+\frac{\epsilon_{s}}{\gamma\sigma^{2}}e^{-r(T-s)}}\in(0,1]$. Then the derivative of social welfare with respect to the cost is
@@
Condition (4.1) restricts that $\epsilon_{s}\in(\underline{\epsilon}_{s},\overline{\epsilon}_{s}]$ if $\underline{\theta}>0$. And $\epsilon_{s}>\underline{\epsilon}_{s}$, $\epsilon_{s}\geq\overline{\epsilon}_{s}$ if $\underline{\theta}<0$. As we assume that $\overline{\epsilon}_{s}<0$, and $\underline{\epsilon}_{s}<0$, feasible $\epsilon_{s}$ ranges from 00 to $\overline{\epsilon}$. If $\rho<0$ or $\rho\geq 2\overline{\rho}_{s}$, the welfare always decreases in $\epsilon_{s}$, therefore the optimal cost is $\epsilon^{\ast}_{s}=0$. If $0<\rho<\underline{\rho}_{s}$, the welfare always increases in $\epsilon_{s}$, then the optimal cost is $\epsilon^{\ast}_{s}=\overline{\epsilon}$. If $\underline{\rho}_{s}\leq\rho<2\overline{\rho}_{s}$, the welfare decreases in $\epsilon_{s}$ first and them increases in it, reaching the maximum at 00 or $\overline{\epsilon}$. $\square$
@@ 6 Conclusion
In this paper, we develop a dynamic model to study the general equilibrium of the insurance market, emphasizing the interaction between insurers’ underwriting and investment strategies. Depending on the parameter conditions, three distinct equilibrium outcomes can emerge: positive insurance market, zero insurance market, and market failure. Specifically, insurers may rationally accept underwriting losses while relying on investment returns, particularly when there is a negative correlation between insurance gains and financial returns. This provides a theoretical explanation for the observed behavior of Chinese P&C insurers.

### v2

@@ 2.3 Optimization Objective
Definition 1.
For each $t\in[0,T]$, $a\triangleq\{a_{s}=(x_{s},y_{s}):t\leq s\leq T\}$ is an admissible strategy for the insurer with initial wealth $m_{t}=m$ and denoted as $a\in\mathcal{A}_{t}$ if it satisfies:
@@
    For any $(t,m)\in[0,T]\times\mathbb{R}$, $\mathbb{E}_{t,m}[\sup\limits_{t\leq s\leq T}\left|m_{s}\right|^{2}]<\infty$;
@@
$V(t,m)\triangleq\sup_{a\in\mathcal{A}_{t}}\mathbb{E}_{t,x}[U(m(T))],$ (2.2)
where $V(\cdot,\cdot)$ denotes the value function, with the boundary condition $V(T,m)=U(m)$. Based on this optimization problem, we now provide the definition of the equilibrium of the insurance market.
Definition 2.
A stationary Markovian competitive equilibrium consists of: an insurance price process $\Theta$, an insurance demand process $D$, an insurer’s wealth process $M$, an insurance supply process $X$, and an insurer’s investment process $Y$, that are compatible with the insurer’s optimization problem (2.2) and the market clearing condition $D=X$.
@@ 3.1 Solution to Decision Problem
$V_{t}+\sup_{x\geq 0,y}\left\{V_{m}\left(x\theta l+y(\mu-r)+mr\right)+\frac{1}{2}V_{mm}(x^{2}\eta^{2}+2\rho x\eta y\sigma+y^{2}\sigma^{2})\right\}=0.$ (3.1)
@@
For $\lambda^{\ast}\geq 0$, it requires that $\rho\geq\frac{\theta l\sigma}{(\mu-r)\eta}$. Substituting this result back into the HJB equation (3.1), we obtain
@@
For $x^{\ast}>0$, we require $\rho<\frac{\theta l\sigma}{(\mu-r)\eta}$. Replacing them back into the HJB equation (3.1) gives
@@
Proposition 1.
Define $f(t,m,x,y)=V_{m}\left(x\theta l+y(\mu-r)+mr\right)+\frac{1}{2}V_{mm}\left(x^{2}\eta^{2}+2\rho x\eta y\sigma+y^{2}\sigma^{2}\right)$. Assume $V_{m}>0$ and $V_{mm}<0$ and that $(x^{\ast},y^{\ast})$ is defined as follows:
@@
Then $f(t,m,x^{\ast},y^{\ast})=\sup\limits_{x\geq 0,y}f(t,m,x,y)$.
Consequently, if $a^{\ast}=\left\{a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):(x^{\ast}_{s},y^{\ast}_{s})=(x^{\ast}(t,m^{\ast}_{s}),y^{\ast}(t,m^{\ast}_{s}):t\leq s\leq T\right\}\in\mathcal{A}_{t}$ and the HJB equation (3.1) admits a solution $V$, then $a^{\ast}$ is the optimal control of Problem (2.2), where $M^{\ast}=\{m^{\ast}_{s}:s\geq t\}$ is the unique solution of SDE (2.1) when the process $a$ is replaced with $a^{\ast}$.
@@ 3.2 Case of CARA Utility
Proposition 2.
Given the price process $\Theta$, for an insurer with CARA utility and initial wealth $m$ at time $t$, the optimal underwriting and investment strategy is $a^{\ast}=\left\{a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):t\leq s\leq T\right\}$, and the corresponding value function $V(t,m)$ is
@@
Proof..
Following Yang and Zhang, (2005), we guess that the value function takes the form of (3.2). It can be efficiently verified that $a^{\ast}$ is admissible and $V$ satisfies the HJB equation (3.1). Therefore, they solve Problem (2.2). ∎
Note that now the price $\theta_{s}$ is assumed to be given, and each insurer in the competitive market acts as a price taker. Although the equilibrium price would be eventually determined through the interaction of market supply and demand, the supply function $x^{\ast}_{s}(\theta_{s})$, as shown in (1) and (3.4), is deterministic at each point in time, and the demand function $d(\theta_{s})$ is also deterministic. As a result, the equilibrium price is not stochastic and varies only with time $s$. Insurers hold deterministic beliefs about this equilibrium price process.
@@ 3.3 Market Equilibrium
Now we analyze the equilibrium conditions in the insurance market. In equilibrium, the competitive price of insurance must clear the market, meaning that the aggregate supply of insurance meets the aggregate demand, i.e., the equilibrium price $\theta^{\ast}_{t}$ should satisfy the relationship(or market equilibrium equation):
@@
Theorem 1 (Market Equilibrium).
@@
    $\Theta^{\ast}_{z}=\left\{\theta^{\ast}_{s}=\overline{\theta}:t\leq s\leq T\right\}.$
@@
    $X^{\ast}_{z}(\Theta^{\ast})=\left\{x^{\ast}_{s}(\theta^{\ast}_{s})=0:t\leq s\leq T\right\},$
@@
    then there exists a stopping time $\tau_{f}=\inf\limits_{s\geq t}\{\underline{\theta}>\gamma(1-\rho^{2})\frac{\eta^{2}}{l}e^{r(T-s)}+\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}\}$. When $\tau_{f}\leq s\leq T$, the market fails. Thus, there is no equilibrium in the insurance market.
Proof..
Based on Proposition 2, when $\phi_{s}>\rho$, the insurer actively participates in both underwriting and investment. In this case, the supply function is given by $x^{\ast}_{s}(\theta_{s})=\frac{\theta_{s}l\sigma-\rho(\mu-r)\eta}{\gamma(1-\rho^{2})\eta^{2}\sigma}e^{-r(T-s)}$, which is an upward-sloping linear curve and has an intersection with the downward-sloping demand curve. But to further guarantee the intersection is an equilibrium, we need
@@
When $\phi_{s}\leq\rho$, the insurer stops underwriting, and $x^{\ast}_{s}=0$. The equilibrium price becomes $\theta^{\ast}_{s}=\overline{\theta}$, and it should satisfy $\phi_{s}(\theta^{\ast}_{s})\leq\rho$. Then Condition (3.6) is derived. Beyond these two cases, when $\underline{\theta}$ is set relatively high, the intersection would violate $\underline{\theta}\leq\theta^{\ast}_{s}$, meaning there is no feasible equilibrium. This results in the condition specified in (3.7). ∎
@@ 3.5 Case of Negative Loading
Proposition 5.
Given $\frac{\rho}{1-\rho^{2}}<-\gamma\frac{\sigma}{\mu-r}\frac{\eta\overline{\theta}}{\Delta\theta}$, and the positive insurance market condition (3.5), there exists a stopping time $\tau_{n}=\inf\limits_{s\geq t}\left\{\frac{\rho}{1-\rho^{2}}<-\gamma\frac{\sigma}{\mu-r}\frac{\eta\overline{\theta}}{\Delta\theta}e^{r(T-s)}\right\}$. When $\tau_{n}\leq s\leq T$, the equilibrium loading premium $\theta^{\ast}_{s}$ becomes negative.
@@ 3.6 Welfare Implication
Theorem 2 (Social Welfare).
@@ 3.7 Case of Perfect Correlation
$\Theta^{\ast}=\{\theta^{\ast}_{s}=\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}:t\leq s\leq T\}.$
@@
$X^{\ast}(\Theta^{\ast})=\{x^{\ast}_{s}(\theta^{\ast}_{s})=\frac{\overline{\theta}-\rho\frac{\mu-r}{\sigma}\frac{\eta}{l}}{\Delta\theta}:t\leq s\leq T\},$
@@
$Y^{\ast}(\Theta^{\ast})=\{y^{\ast}_{s}(\theta^{\ast}_{s})=\frac{\frac{\eta^{2}}{\sigma}\left(\frac{\mu-r}{\sigma}-\rho\frac{l}{\eta}\overline{\theta}\right)+\frac{\mu-r}{\gamma\sigma^{2}}l\Delta\theta e^{-r(T-s)}}{l\Delta\theta}:t\leq s\leq T\}.$
@@ 4.1 Market Equilibrium
$\mathrm{d}m_{t}=\big{(}x_{t}\theta_{t}l+y_{t}(\mu-r)-\frac{1}{2}y_{t}^{2}\epsilon_{t}+m_{t}r\big{)}\mathrm{d}t+x_{t}\eta\mathrm{d}W^{I}_{t}+y_{t}\sigma\mathrm{d}W^{S}_{t},$
where $\epsilon_{t}\in[0,\overline{\epsilon}]$ denotes the intensity of regulatory frictions and consists of $\mathcal{E}\triangleq\left\{\epsilon_{t}:t\geq 0\right\}$. And $\overline{\epsilon}$ is an exogenously given bar. Using a similar approach to the derivation in Section 3, we obtain the following proposition describing the insurer’s optimal strategy.
Proposition 7.
Given the price process $\Theta$ and the regulatory cost process $\mathcal{E}$, for an insurer with CARA utility and initial wealth $m$ at time $t$, the optimal underwriting and investment strategy is $a^{\ast}=\left\{a^{\ast}_{s}=(x^{\ast}_{s},y^{\ast}_{s}):t\leq s\leq T\right\}$, and corresponding value function $V(t,m)$ is
@@
Theorem 3 (Market Equilibrium with Regulatory Friction).
@@ 4.2 Welfare Implication
In this subsection, we examine how regulatory frictions impact the social welfare of the insurance market and explore the conditions for the regulator to set an optimal cost.
Theorem 4 (Social Welfare with Regulatory Friction).
@@
Proof..
The social welfare can be easily calculated by $w_{s}^{\ast}=\frac{1}{2}x_{s}^{\ast}\left(\overline{\theta}-\frac{\rho(\mu-r)\eta}{\left(1+\frac{\epsilon_{s}}{\gamma\sigma^{2}}e^{-r(T-s)}\right)l\sigma}\right)$. Define $\alpha_{s}=\frac{1}{1+\frac{\epsilon_{s}}{\gamma\sigma^{2}}e^{-r(T-s)}}\in(0,1]$. Then the derivative of social welfare with respect to the cost is
@@
Condition (4.1) restricts that $\epsilon_{s}\in(\underline{\epsilon}_{s},\overline{\epsilon}_{s}]$ if $\underline{\theta}>0$. And $\epsilon_{s}>\underline{\epsilon}_{s}$, $\epsilon_{s}\geq\overline{\epsilon}_{s}$ if $\underline{\theta}<0$. As we assume that $\overline{\epsilon}_{s}<0$, and $\underline{\epsilon}_{s}<0$, feasible $\epsilon_{s}$ ranges from 00 to $\overline{\epsilon}$. If $\rho<0$ or $\rho\geq 2\overline{\rho}_{s}$, the welfare always decreases in $\epsilon_{s}$, therefore the optimal cost is $\epsilon^{\ast}_{s}=0$. If $0<\rho<\underline{\rho}_{s}$, the welfare always increases in $\epsilon_{s}$, then the optimal cost is $\epsilon^{\ast}_{s}=\overline{\epsilon}$. If $\underline{\rho}_{s}\leq\rho<2\overline{\rho}_{s}$, the welfare decreases in $\epsilon_{s}$ first and them increases in it, reaching the maximum at 00 or $\overline{\epsilon}$. ∎
@@ 6 Conclusion
In this paper, we develop a dynamic model to study the equilibrium of insurance market, emphasizing the interaction between insurers’ underwriting and investment strategies. Depending on the parameter conditions, three distinct equilibrium outcomes can emerge: positive insurance market, zero insurance market, and market failure. Specifically, insurers may rationally accept underwriting losses while relying on investment returns, particularly when there is a negative correlation between insurance gains and financial returns. This provides a theoretical explanation for the observed behavior of Chinese P&C insurers.